OpenStudy (poetryprincess):

For what values of m does the graph of y = 3x^2 + 7x + m have two x-intercepts? a) m>12/49 b) m<12/49 c) m<49/12 d) m> 49/12

OpenStudy (owlcoffee):

The quadratic formula, has a part we call the "discriminant" defined by the variables that are inside the square root, and is denotated by "delta": \[\Delta = b^2 - 4ac\] Whenever we solve a quadratic equation that is complete and we analyze the discriminant, we can get 3 scenarios: \[if \rightarrow \Delta > 0 =>\exists x_1, x_2 / ax^2 +bx+c=0\] This just means: "if the discriminant is greater than zero, there will exist two x-intercepts" And for the second scenario: \[if \rightarrow \Delta = 0 \rightarrow \exists x_o / ax^2 + bx + c=0\] This means: "if the discriminant is equal to zero, there will be one and only one x-intercept" And for the last scenario: \[if \rightarrow \Delta < 0 \rightarrow \exists x \notin \mathbb{R} / ax^2 + bx + c=0\] This means that :"if the discriminant is less than zero, there will be no x-intercepts" So, if we take your excercise and analyze the the discriminant: \[3x^2 + 7x+m = y\] we will find the values that satisfy y=0 : \[3x^2 + 7x +m =0\] And we'll analyze the discriminant: \[\Delta=7^2 - 4(3)(m)\] And we are only interested in the values that make the discriminant equal zero: \[7^2 - 4(3)(m)=0\] All you have to do is solve for "m".